# Parametrically-controlled microwave-photonic interface for the fluxonium

### v2

@@ I Introduction
In recent years, the fluxonium [20] has emerged as a promising superconducting qubit due to its remarkably long coherence times [21, 22] and high gate fidelities [23, 24]. A coherent fluxonium-photonic interface would be beneficial for fluxonium quantum processing on multiple fronts. First, efficient initialization of the fluxonium is important as the low transition frequency of the fluxonium results in a highly-mixed thermal population. Second, faithful mapping between a fluxonium excitation and a flying qubit would allow the realization of modular quantum systems linked via traveling microwave photons using the fluxonium. A typical method for engineering this interface involves converting a fluxonium excitation into a fast-decaying resonator photon through pumping of cavity sideband transitions [25, 26, 27]. Recent experiments have reported multiple ways of circumventing the parity selection rule forbidding these direct transitions by flux-biasing the fluxonium qubit away from half flux [25], using microwave-driven multi-photon transitions [26], and coupling the fluxonium to a resonator via a third-order mixing element known as the Superconducting Nonlinear Asymmetric Inductive eLement (SNAIL) [28, 27]. The above-mentioned experiments have been solely focused on fluxonium initialization and cooling.
@@ II Fluxonium-SNAIL-resonator system
$\hat{H}_{\text{sys}}/\hbar=\sum_{i=f,s,r}4E_{C,i}\hat{n}_{i}^{2}+\frac{1}{2}E_{L,i}\hat{\phi}_{i}^{2}-E_{J,f}\cos{(\hat{\phi}_{f}-\phi_{\text{ext}}^{f})}\\ +c_{3}^{s}\hat{\phi}_{s}^{3}+g_{fs}\hat{\phi}_{f}\hat{\phi}_{s}+g_{sr}\hat{n}_{s}\hat{n}_{r}.$ (1)
@@
with $|\xi(t)|^{2}$ being the number of pump photons in the SNAIL resonator, and $\phi_{i,\mathrm{zpf}}$ denoting the zero-point fluctuations of the flux associated with each mode (Appendix C). This Hamiltonian allows us to drive transitions of the form $|i,n\rangle\leftrightarrow|j,n\pm 1\rangle,~{}i\neq j$ and $n>0$, where $i,j$ denote any neighbouring eigenstates of the fluxonium and $n$ denotes the number of photons in the readout resonator. By directly driving the $|e0\rangle\leftrightarrow|g1\rangle$ transition, excitations of the fluxonium can be parametrically converted into a photon in the readout resonator, which is then emitted at the decay rate $\kappa$ (Fig. 1(b)).
The rate of conversion is controlled by the pump amplitude $\xi(t)$ and the third-order coefficient $c_{3}^{s}$. A strong microwave drive on the SNAIL induces a significant ac-Stark shift on the transition (see Appendix C for further details on the system Hamiltonian), which complicates pulse calibration due to the Stark shift changing non-negligibly during the pulse rise time. We therefore choose to work with a large $c_{3}^{s}$, which can be obtained by operating at nonzero $\Phi_{\text{ext}}^{s}$ through the SNAIL. We flux bias the system to $\Phi_{\text{ext}}^{f}/\Phi_{\text{0}}=9.5$, corresponding to $\Phi_{\text{ext}}^{s}/\Phi_{\text{0}}=0.27$ (Appendix A). Spectroscopy of the desired $|e0\rangle\leftrightarrow|g1\rangle$ transition at this flux point is shown in Fig. 1(c). We note the clear visibility of this transition at the half-flux sweet spot of the fluxonium, indicating that the parity selection rule has been broken [27]. The expected $c^{s}_{3}/2\pi$ is simulated to be 350 MHz at this point.
@@ III Coherent Driving
Here we demonstrate that our parametric scheme can reach different coupling regimes useful for fast reset and generating flying photons. We characterize the transition $|e0\rangle\leftrightarrow|g1\rangle$ by driving it with a square pulse of varying duration and power (Fig. 2). We measure the qubit population in time as a function of detuning $\omega_{d}-\omega_{0}$, where $\omega_{d},\omega_{0}$ represent the driving frequency and the transition frequency, respectively (Fig. 2(a)). As the system evolves coherently as a superposition of $|e0\rangle$ and $|g1\rangle$, $|g1\rangle$ decays back to $|g0\rangle$ and emits the photon field into the transmission line at an effective decay rate depending on the relative strength of $g$ and $\kappa$. Before readout, we wait for time $t_{\text{buffer}}=5$ $\mu$s ($\sim 10\kappa^{-1}$), to ensure that the photon has left the resonator to avoid changes to the fluxonium readout based on the cavity state.
@@
We then characterize the $|e0\rangle\leftrightarrow|g1\rangle$ transition in the weak coupling regime, which is useful for efficient generation and temporal shaping of flying photons. Under a relatively weak pump, the state $|e0\rangle$ is slowly converted into $|g1\rangle$ with linewidth $\kappa$. In this regime, we can tune the effective decay rate $\Gamma_{\mathrm{eff}}=4g^{2}/\kappa$ via the strength of our pump, which controls $g$. In Fig. 2(d), we show the qubit $|e0\rangle$ population evolution in time as a function of detuning $\omega_{d}-\omega_{0}$ when the $|e0\rangle\leftrightarrow|g1\rangle$ transition is driven by a relatively weak pump tone. The data at zero detuning is plotted separately in Fig. 2(e). A fit to the master equation (Appendix C) yields a coupling rate $g/2\pi=0.14\text{ MHz}\ll\kappa/2\pi$. The qubit decay exhibits no oscillations and resembles that of standard Purcell loss after the ring-down time $1/\kappa$ with effective decay rate $\Gamma_{\mathrm{eff}}=4g^{2}/\kappa$. Now with a well-characterized ability to drive and tune the $|e0\rangle\leftrightarrow|g1\rangle$ transition, we can use this transition for fast reset and controlled release of single photons.
@@ IV Fast Reset
Here we perform fast reset of the fluxonium in the strong coupling regime. Specifically, we turn on a strong pump for a pulse duration $t=200$ ns for fast evacuation of the $|e0\rangle$ population into the $|g1\rangle$ state. We then turn off the drive and wait for $t_{\mathrm{buffer}}=5\ \mu$s to allow the photon to leave before readout. In Fig. 3(a), we show the measured fluxonium population in $|e0\rangle$ as a function of the reset pulse duration for a fixed pulse amplitude. As we increase the pulse duration, the fluxonium population in $|e0\rangle$ decreases and reaches the maximum population in $|g0\rangle$ with pulse duration $t=200~{}$ns before the excitation gets swapped back to $|e0\rangle$. At thermal equilibrium, the qubit ($\omega_{f}/2\pi=81$ MHz) has a probability of $55\%$ to be in the ground state, which is equivalent to a qubit temperature of 20 mK (Fig. 3(b)). Using our reset protocol, we measure a maximum of $92\%$ qubit population in the ground state, which corresponds to cooling the qubit to 1.6 mK (Fig. 3(c)). Both the long timescale and the observed inefficiency of the reset is due to the low decay rate $\kappa/2\pi=0.4$ MHz of the resonator. The qubit heats during $t_{\text{buffer}}$ and readout. If we correct for the readout error induced due to the qubit population change, the reset efficiency rises to $\sim 96\%$.
@@ V Controlled Release
For characterization of the controlled conversion process, we analyze the emitted photon superposition state of the $(|0\rangle+e^{i\phi}|1\rangle)/\sqrt{2}$, which has a non-zero field amplitude. To initialize the qubit in states $|\psi\rangle=(|g0\rangle+e^{i\phi}|e0\rangle)/\sqrt{2}$, we apply a reset pulse followed by a $\pi/2$ pulse of phase $\phi$ on the qubit. A traveling photon of state $(|0\rangle+e^{i\phi}|1\rangle)/\sqrt{2}$ is then controllably released by pumping the fluxonium at the $|e0\rangle\rightarrow|g1\rangle$ transition frequency in the regime where $g\ll\kappa$ (pulse sequence shown Fig. 4(a)). The outgoing field is recorded using heterodyne detection, which also captures the phase information of the emitted photon.
For initial characterization, we perform this controlled release protocol using a simple square pulse. In Fig. 4(b), we show the measured in-phase quadrature of the propagating fields as a function of time, with the fluxonium initialized to be in $|\psi^{\pm}\rangle=(|g0\rangle\pm|e0\rangle)/\sqrt{2}$. With a square pulse, the coupling rate of $|e0\rangle\leftrightarrow|g1\rangle$ is constant, which should result in a simple exponential damping of the output fields after some ring-up time proportional to $\kappa^{-1}$ as shown in Fig. 4(b). In this case, the exact response of the resonator as well as the emitted photon related by $\hat{r}_{\text{out}}=\sqrt{\kappa}\hat{r}$ can be analytically computed using input-output theory (Appendix C), which fits our measured data well. Here the coupling rate $g(t)$ is fit to be $2\pi\times 0.1\ \mathrm{MHz}\ll\kappa$. Upon shifting the phase of the qubit Rabi pulse by $\pi$, the response of the in-phase signal also shifts correspondingly, which indicates the phase coherence of the controlled release process.
@@
where $\gamma_{\text{ph}}$ is the photon bandwidth (Appendix E). In Fig. 4(c), we show the measured in-phase quadrature and field amplitude of the shaped wave packet with the fluxonium initialized to be in $|\psi^{+}\rangle$, which displays a clear hyperbolic secant shape. This emitted wavepacket could then be captured by a receiver by time-reversing the emission process [2].
@@ VI Outlook
We have demonstrated parametric conversion between a fluxonium qubit excitation and a flying photon by successfully driving the $|e0\rangle\leftrightarrow|g1\rangle$ transition. Utilizing this transition, we have initialized the fluxonium qubit in the ground state with measured probability $>95\%$ and predict a reset threshold in the same system of $>99\%$ with a longer reset pulse. By using a resonator with a higher decay rate, our simulations indicate that the fluxonium could be reset with fidelity $\sim 99.8\%$ in 500 ns. We have also demonstrated coherent release and temporal shaping of a flying photon detected using heterodyne measurement. By introducing a microwave-photonic interface for the fluxonium qubit, we have paved the way for remotely entangling fluxonium qubits via a microwave-photonic link.
@@ Appendix A Device characterization
Figure 5 shows the optical and scanning electron microscope (SEM) images of our device. To drive the parity-forbidden transition $|e0\rangle\leftrightarrow|g1\rangle$, we need strong coupling between the fluxonium and the SNAIL, and between the SNAIL and the readout resonator respectively. We realize strong coupling by using multiple SNAILs as a shared inductance with the fluxonium superinductance. Moreover, one side of the SNAIL capacitor is enlarged for stronger coupling capacitance with the readout resonator and asymmetric driving of the system. A charge line weakly coupled to the other pad of the SNAIL resonator is included for driving. The coupling between the charge line and the SNAIL is designed as a trade off between the need to strongly drive the parity-forbidden transitions and while still suppressing the Purcell loss of the fluxonium excitation through the SNAIL to the charge line. The readout resonator is coupled to the readout line in the hanger style with relatively low decay rate $\kappa/2\pi=0.4$ MHz. We fabricate the entire device, with the exception of Al/AlOx/Al junctions, with tantalum on a sapphire wafer. A global coil is used to tune the flux across the entire chip, including flux threading though the fluxonium $\Phi_{\text{ext}}^{f}$ and the SNAIL $\Phi_{\text{ext}}^{s}$. This is made possible by designing a large area ratio between the fluxonium and the SNAIL $A_{f}/A_{s}\approx 35$, so that one may tune either flux independently. We operate our device at $\Phi_{\text{ext}}^{f}/\Phi_{0}=9.5$, corresponding to $\Phi_{\text{ext}}^{s}/\Phi_{\text{0}}=0.27$. A summary of our device’s parameters at $\Phi_{\text{ext}}^{f}/\Phi_{0}=9.5$ are provided in Table I. Two-tone spectroscopy data for the fluxonium qubit is shown in Fig. 6(a).

### v3

@@ I Introduction
In recent years, the fluxonium [20] has emerged as a promising superconducting qubit due to its remarkably long coherence times [21, 22] and high gate fidelities [23, 24]. A coherent fluxonium-photonic interface would be beneficial for fluxonium quantum processing on multiple fronts. First, efficient initialization of the fluxonium is important as the low transition frequency of the fluxonium results in a highly-mixed thermal population. Second, faithful mapping between a fluxonium excitation and a flying qubit would allow the realization of modular quantum systems linked via traveling microwave photons using the fluxonium. A typical method for engineering this interface involves converting a fluxonium excitation into a fast-decaying resonator photon through pumping of cavity sideband transitions [25, 26, 27]. However, direct sideband transitions between fluxonium-resonator states of the same total parity are forbidden under a linear microwave drive. Recent experiments have reported multiple ways of circumventing the parity selection rule forbidding these direct transitions by flux-biasing the fluxonium qubit away from half flux [25], using microwave-driven multi-photon transitions [26], and coupling the fluxonium to a resonator via a third-order mixing element known as the Superconducting Nonlinear Asymmetric Inductive eLement (SNAIL) [28, 27]. The above-mentioned experiments have been solely focused on fluxonium initialization and cooling.
@@ II Fluxonium-SNAIL-resonator system
$\hat{H}_{\text{sys}}/\hbar=\sum_{i=f,s,r}(4E_{C,i}\hat{n}_{i}^{2}+\frac{1}{2}E_{L,i}\hat{\phi}_{i}^{2})\\ -E_{J,f}\cos{(\hat{\phi}_{f}-\phi_{\text{ext}}^{f})}+c_{3}^{s}\hat{\phi}_{s}^{3}+g_{fs}\hat{\phi}_{f}\hat{\phi}_{s}+g_{sr}\hat{n}_{s}\hat{n}_{r}.$ (1)
@@
with $|\xi(t)|^{2}$ being the number of pump photons in the SNAIL resonator, and $\phi_{i,\mathrm{zpf}}$ denoting the zero-point fluctuations of the flux associated with each mode (Appendix C). This Hamiltonian allows us to drive transitions of the form $|i,n\rangle\leftrightarrow|j,n+1\rangle,~{}i\neq j$ and $n>0$, where $i,j$ denote any neighbouring eigenstates of the fluxonium with $|i\rangle$ being the higher eigenstate, and $n$ denotes the number of photons in the readout resonator. By directly driving the $|e0\rangle\leftrightarrow|g1\rangle$ transition, excitations of the fluxonium can be parametrically converted into a photon in the readout resonator, which is then emitted at the decay rate $\kappa$ (Fig. 1(b)).
The rate of conversion is controlled by the pump amplitude $\xi(t)$ and the third-order coefficient $c_{3}^{s}$. A strong microwave drive on the SNAIL induces a significant ac-Stark shift on the transition (see Appendix C for further details on the system Hamiltonian and ac-Stark shift), which complicates pulse calibration due to the Stark shift changing non-negligibly during the pulse rise time. We therefore choose to work with a large $c_{3}^{s}$, which can be obtained by operating at nonzero $\Phi_{\text{ext}}^{s}$ through the SNAIL. We flux bias the system to $\Phi_{\text{ext}}^{f}/\Phi_{\text{0}}=9.5$, corresponding to $\Phi_{\text{ext}}^{s}/\Phi_{\text{0}}=0.27$ (Appendix A). Spectroscopy of the desired $|e0\rangle\leftrightarrow|g1\rangle$ transition at this flux point is shown in Fig. 1(c) (see Appendix B for wiring). We perform dispersive readout on the fluxonium enabled by the last two, always-on coupling terms in Eq. (1). We note the clear visibility of this transition at the half-flux sweet spot of the fluxonium, indicating that the parity selection rule has been broken [27]. The expected $c^{s}_{3}/2\pi$ is simulated to be 350 MHz at this point.
@@ III Coherent Driving
Here we demonstrate that our parametric scheme can reach different coupling regimes useful for fast reset and generating flying photons. We characterize the transition $|e0\rangle\leftrightarrow|g1\rangle$ by driving it with a square pulse of varying duration and power (Fig. 2). We measure the qubit population in time as a function of detuning $\omega_{d}-\omega_{0}$, where $\omega_{d},\omega_{0}$ represent the driving frequency and the transition frequency, respectively (Fig. 2(a)). As the system evolves coherently as a superposition of $|e0\rangle$ and $|g1\rangle$, $|g1\rangle$ decays back to $|g0\rangle$ and emits the photon field into the transmission line at an effective decay rate depending on the relative strength of $g$ and $\kappa$. We then wait for time $t_{\text{buffer}}=5$ $\mu$s ($\sim 10\kappa^{-1}$), to ensure that the photon has left the resonator to avoid changes to the qubit readout as well as frequency shifts based on the cavity state before any subsequent operation on the fluxonium.
@@
We then characterize the $|e0\rangle\leftrightarrow|g1\rangle$ transition in the weak coupling regime, which is useful for efficient generation and temporal shaping of flying photons. Under a relatively weak pump, the state $|e0\rangle$ is slowly converted into $|g1\rangle$ with linewidth $\kappa$. In this regime, we can tune the effective decay rate $\Gamma_{\mathrm{eff}}=4g^{2}/\kappa$ via the strength of our pump, which controls $g$. In Fig. 2(d), we show the qubit $|e0\rangle$ population evolution in time as a function of detuning $\omega_{d}-\omega_{0}$ when the $|e0\rangle\leftrightarrow|g1\rangle$ transition is driven by a relatively weak pump tone. The data at zero detuning is plotted separately in Fig. 2(e). A fit to the master equation (Appendix C) yields a coupling rate $g/2\pi=0.13\text{ MHz}\ll\kappa/2\pi$. The qubit decay exhibits no oscillations and resembles that of standard Purcell loss after the ring-down time $1/\kappa$ with effective decay rate $\Gamma_{\mathrm{eff}}=4g^{2}/\kappa$. Now with a well-characterized ability to drive and tune the $|e0\rangle\leftrightarrow|g1\rangle$ transition, we can use this transition for fast reset and controlled release of single photons.
@@ IV Fast Reset
Here we perform fast reset of the fluxonium in the strong coupling regime. Specifically, we turn on a strong pump for a pulse duration $t=200$ ns for fast evacuation of the $|e0\rangle$ population into the $|g1\rangle$ state. We then turn off the drive and wait for $t_{\mathrm{buffer}}=5\ \mu$s to allow the photon to leave before readout. In Fig. 3(a), we show the measured fluxonium population in $|e0\rangle$ as a function of the reset pulse duration for a fixed pulse amplitude. As we increase the pulse duration, the fluxonium population in $|e0\rangle$ decreases and reaches the maximum population in $|g0\rangle$ with pulse duration $t=200~{}$ns before the excitation gets swapped back to $|g0\rangle$. At thermal equilibrium, the qubit ($\omega_{f}/2\pi=81$ MHz) has a probability of $55\%$ to be in the ground state, which is equivalent to a qubit temperature of 20 mK (Fig. 3(b)). Using our reset protocol, we measure a maximum of $92\%$ qubit population in the ground state, which corresponds to cooling the qubit to 1.6 mK (Fig. 3(c)). Both the long timescale and the observed inefficiency of the reset is due to the low decay rate $\kappa/2\pi=0.4$ MHz of the resonator. The qubit heats during $t_{\text{buffer}}$ and readout. If we correct for the readout error induced due to the qubit population change, the reset efficiency rises to $\sim 96\%$.
@@ V Controlled Release
For characterization of the controlled conversion process, we analyze the emitted photon superposition state of the $(|0\rangle+e^{i\phi}|1\rangle)/\sqrt{2}$, which has a non-zero field amplitude. To initialize the qubit in states $|\psi\rangle=(|g0\rangle+e^{i\phi}|e0\rangle)/\sqrt{2}$, we apply a reset pulse followed by a $\pi/2$ pulse of phase $\phi$ on the qubit. A traveling photon of state $(|0\rangle+e^{i\phi}|1\rangle)/\sqrt{2}$ is then controllably released by pumping the fluxonium at the $|e0\rangle\rightarrow|g1\rangle$ transition frequency in the regime where $g\ll\kappa$ (pulse sequence shown in Fig. 4(a)). The outgoing field is recorded using heterodyne detection, which also captures the phase information of the emitted photon.
For initial characterization, we perform this controlled release protocol using a simple square pulse. In Fig. 4(b), we show the measured in-phase quadrature of the propagating fields as a function of time, with the fluxonium initialized to be in $|\psi^{\pm}\rangle=(|g0\rangle\pm|e0\rangle)/\sqrt{2}$. With a square pulse, the coupling rate of $|e0\rangle\leftrightarrow|g1\rangle$ is constant, which should result in a simple exponential damping of the output fields with $\Gamma_{\mathrm{eff}}=4g^{2}/\kappa$ after a ring-up time proportional to $\kappa^{-1}$ as shown in Fig. 4(b). In this case, the exact response of the resonator as well as the emitted photon related by $\hat{r}_{\text{out}}=\sqrt{\kappa}\hat{r}$ can be analytically computed using input-output theory (Appendix C), which fits our measured data well. Here the coupling rate $g(t)$ is fit to be $2\pi\times 0.1\ \mathrm{MHz}\ll\kappa$. Upon shifting the phase of the qubit Rabi pulse by $\pi$, the response of the in-phase signal also shifts correspondingly, which indicates the phase coherence of the controlled release process.
@@
where $\gamma_{\text{ph}}$ is the photon bandwidth. In Fig. 4(c), we show the measured in-phase quadrature and field amplitude of the shaped wave packet with the fluxonium initialized to be in $|\psi^{+}\rangle$, which displays a clear hyperbolic secant shape. Based on numerical simulation, we estimate a conversion efficiency of $\eta\approx 79\%$ taking into account state preparation errors, which is mainly due to qubit depolarization during the release process. By using a resonator of higher decay rate $\kappa/2\pi=4\mathrm{\ MHz}$, the conversion process can be shortened down to $\lesssim 1\ \mu$s (see Fig. 12 in Appendix E). In this case, conversion loss due to qubit depolarization can be easily suppressed with no additional system optimization and we simulate a conversion efficiency of $\eta\approx 99\%$ assuming our current qubit lifetimes and perfect state initialization. We quantify the symmetry of the wavepacket by computing its normalized scalar product with the time-reversed copy of itself [17], which yields $s=0.975$ (Eq. (31)). This emitted wavepacket could then be captured by a receiver by time-reversing the emission process [2].
@@ VI Outlook
We have demonstrated parametric conversion between a fluxonium qubit excitation and a flying photon by successfully driving the $|e0\rangle\leftrightarrow|g1\rangle$ transition. Utilizing this transition, we have initialized the fluxonium qubit in the ground state with measured probability $>95\%$ and predict a reset threshold in the same system of $>99\%$ with a longer reset pulse. By using a resonator with a higher decay rate, our simulations indicate that the fluxonium could be reset with fidelity $\sim 99.8\%$ in 500 ns. We have also demonstrated coherent release and temporal shaping of a flying photon detected using heterodyne measurement with total efficiency of $\sim 79\%$. An improved conversion efficiency of $\sim 99\%$ can be straightforwardedly obtained by increasing the resonator decay rate. By introducing a microwave-photonic interface for the fluxonium qubit, we have paved the way for remotely entangling fluxonium qubits via a microwave-photonic link.
@@ Appendix A Device characterization
Figure 5 shows the optical and scanning electron microscope (SEM) images of our device. To drive the parity-forbidden transition $|e0\rangle\leftrightarrow|g1\rangle$, we need strong coupling between the fluxonium and the SNAIL, and between the SNAIL and the readout resonator respectively. We realize strong coupling by using multiple SNAILs as a shared inductance with the fluxonium superinductance. Moreover, one side of the SNAIL capacitor is enlarged for stronger coupling capacitance with the readout resonator and asymmetric driving of the system. A charge line weakly coupled to the other pad of the SNAIL resonator is included for driving. The coupling between the charge line and the SNAIL is designed as a trade off between the need to strongly drive the parity-forbidden transitions and while still suppressing the Purcell loss of the fluxonium excitation through the SNAIL to the charge line. The readout resonator is coupled to the readout line in hanger style with relatively low decay rate $\kappa/2\pi=0.4$ MHz. Single-tone spectroscopy for the readout resonator and two-tone spectroscopy for the fluxonium qubit is shown in Fig. 6 and Fig. 7(a) respectively. The resonator transmission coefficient $S_{21}(f)$ in the complex plane [29] is fit to a standard hanger geometry response:
$S_{21}(f)=ae^{i\alpha}(1-\frac{Q/Q_{c}}{1+2iQ\frac{f-f_{0}}{f_{0}}}),$ (5)
where $f_{0}$ is the resonance frequency, $Q$ and $Q_{c}$ denote the total and external quality factor, $a$ attenuates the $S_{21}$ radius while $\alpha$ rotates the $S_{21}$ in the complex plane. The fluxonium spectroscopy is fit by numerically diagonalizing the Hamiltonian in Eq. 1. A summary of our device’s parameters at $\Phi_{\text{ext}}^{f}/\Phi_{0}=9.5$ are provided in Table I.
We fabricate the entire device, with the exception of Al/AlOx/Al junctions, with tantalum on a sapphire wafer. A global coil is used to tune the flux across the entire chip, including flux threading though the fluxonium $\Phi_{\text{ext}}^{f}$ and the SNAIL $\Phi_{\text{ext}}^{s}$. This is made possible by designing a large area ratio between the fluxonium and the SNAIL $A_{f}/A_{s}\approx 35$, so that one may tune either flux independently. We operate our device at $\Phi_{\text{ext}}^{f}/\Phi_{0}=9.5$, corresponding to $\Phi_{\text{ext}}^{s}/\Phi_{\text{0}}=0.27$.
@@
Figure 6: (a) $S_{21}$ response of the readout resonator at single photon level and zero flux plotted in the complex plane. (b) The same $S_{21}$ data plotted in magnitude (top) and phase (bottom).